# Continual Adversarial Defense

### v4

@@ Abstract
In response to the rapidly evolving nature of adversarial attacks against visual classifiers on a monthly basis, numerous defenses have been proposed to generalize against as many known attacks as possible. However, designing a defense method that generalizes to all types of attacks is not realistic because the environment in which defense systems operate is dynamic and comprises various unique attacks that emerge as time goes on. A well-matched approach to the dynamic environment lies in a defense system that continuously collects adversarial data online to quickly improve itself. Therefore, we put forward a practical defense deployment against a challenging threat model and propose, for the first time, the Continual Adversarial Defense (CAD) framework that adapts to attack sequences under four principles: (1) continual adaptation to new attacks without catastrophic forgetting, (2) few-shot adaptation, (3) memory-efficient adaptation, and (4) high accuracy on both clean and adversarial data. We explore and integrate cutting-edge continual learning, few-shot learning, and ensemble learning techniques to qualify the principles. Extensive experiments validate the effectiveness of our approach against multiple stages of modern adversarial attacks and demonstrate significant improvements over numerous baseline methods. In particular, CAD is capable of quickly adapting with minimal budget and a low cost of defense failure while maintaining good performance against previous attacks. Our research sheds light on a brand-new paradigm for continual defense adaptation against dynamic and evolving attacks.  [^1] Code at https://github.com/cc13qq/CAD
@@
Figure 1: In a dynamic environment, the target model $f_{t}$ faces a diverse range of attacks at different stages and is protected by the defender. During the $i$-th stage, the defense model leverages few-shot defense budget provided by the web crawler, security department, and users of the target model to adapt to new attacks. The defense model also needs to efficiently use memory to store the cache while retaining knowledge of past attacks.
Introduction
@@
Adversarial attack (Madry et al. 2017) aims to deceive deep neural networks (DNNs) by adding subtle perturbations to input images, seriously jeopardizing the reliability of DNNs, particularly in domains sensitive to security and trust. In order to promote the robustness of DNN, adversarial training (Zhang et al. 2019) and purification techniques (Yoon, Hwang, and Lee 2021) are proposed to defend against adversarial attacks by one-shot training in which the model enters a static phase after a single defense training stage (Zhou and Hua 2024), but often exhibit reduced classification accuracy in adversarial examples and can sacrifice classification capacity in clean images.
@@
Contrary to the one-shot defense assumption, the defense system deployed in real scenarios encounters a dynamic environment characterized by numerous unique attacks with higher aggressivity (Li et al. 2020). A well-matched approach to the dynamic environment lies in a constantly evolving defense strategy that continuously adapts to newly emerging adversarial algorithms. A concurrent continual defense work (Zhou and Hua 2024) treats attacks as a task sequence and tries to train a robust model through lifelong learning. However, poor performance in both adversarial and clean data hinders its deployment in reality and calls into question the reasonableness of the continual defense setting.
@@
To eliminate the insufficiency of the above defense setting, we put forward a practical defense deployment and a challenging threat model. As depicted in Figure 1, the defense system is deployed in the cloud and continuously adapts to new emerging attacks from stage to stage using defense budget (i.e. adversarial examples and their ground-truths) sourced from the web crawler, security department, and users of the target model in real-world scenarios. In reality, a stage can consist of only a few samples from one or multiple attacks, and a new stage begins once the system gathers enough defense budget. When adapted to new emerging attacks, the defense system is permitted to use defense budget of previous stages cached in memory. The defense system also has access to an initial attack and shares the same training data and architecture as the target model. Meanwhile, attacks are initiated under the gray-box setting, wherein they possess knowledge of the classifier’s architecture and have access to training data, but remain unaware of the defense mechanism implemented.
@@
In this paper, we propose the first Continual Adversarial Defense (CAD) framework, which aims to defend against evolving attacks in a stage-by-stage manner using few-shot defense budget and efficient memory. In the first stage, we use adversarial data from the initial attack to train an initial defense model which is specifically designed to classify adversarial data, complementing the target model. Drawing inspiration from continual learning (CL), we expand the classification layer to adapt to the new attack as incremental classes and then fine-tune this expanded layer using few-shot defense budget. To address the over-fitting issue that arises from the few-shot fine-tuning, we compress the embedding of previous attacks by generating and assigning virtual prototypes to reserve embedding space in the defense model for future attacks. To optimize memory usage in the data domain, we employ prototype augmentation which allows us to maintain the decision boundary of previous stages without the need to store any budget explicitly. After adaptation of one stage, we refine the classifier to prevent undesirable expansion. Simultaneously, a small model is utilized to ensemble the defense model with the target model by estimating reliable logits for input data, ensuring maximum classification accuracy for both clean and adversarial data.
Based on comprehensive experiments conducted on CIFAR-10 and ImageNet-100, the CAD framework demonstrates strong performance in defending against multi-stage attacks relying on few-shot defense budget while maintaining high accuracy on clean data. To the best of our knowledge, this is the first time to discuss the concept and deployment scenario of continual adversarial defense.
@@
$\bullet$ To withstand the rapidly evolving adversarial attacks targeting DNNs, we put forward the deployment of continual defense strategy and propose a practical threat model in which diverse attacks emerge at different stages. The defense system must adapt well to these attacks using a few-shot defense budget and efficient memory management, while also retaining knowledge of past attacks and maintaining high performance on both clean and adversarial data.
$\bullet$ We propose, for the first time, the Continual Adversarial Defense (CAD) framework that defends against attacks in the dynamic scenario under four practical principles: continual adaptation without catastrophic forgetting, few-shot adaptation, memory-efficient adaptation, and high accuracy in classifying both clean and adversarial data.We leverage cutting-edge techniques including continual learning, few-shot learning, non-exemplar class incremental learning, and ensemble learning to satisfy the principles.
@@
Related Work
Adversarial Defense
@@
Researchers have proposed various general defense methods against adversarial attacks (Madry et al. 2017; Wu et al. 2021; Lin et al. 2019; Wang and He 2021). As a standard defense method, adversarial training (Zhang et al. 2020, 2019; Grathwohl et al. 2020; Wang et al. 2023b) aims to enhance the robustness of the target model by training it with adversarial examples. Another branch of adversarial defense involves purifying the data stream (Salman et al. 2020; Hill, Mitchell, and Zhu 2020; Yoon, Hwang, and Lee 2021; Nie et al. 2022) to remove potential adversarial perturbations or noise that could deceive the model. In fact, all of the above methods have shown limitations in robustness, resulting in a lower classification performance on both adversarial and clean data. In this paper, we present the first Continual Adversarial Defense (CAD) framework as a solution to effectively counter a wide spectrum of attacks emerging stage by stage, while ensuring robust performance in classifying clean data.
Continual Learning
@@
Continual learning (CL) (Liu et al. 2022; Kirkpatrick et al. 2017; Wang et al. 2022; Zhou et al. 2022) aims to learn from a sequence of new classes without forgetting old ones and attracts a lot of attention to various computer vision tasks. In recent years, some methods have been aimed at solving the CL problem without relying on preserving data (called no exemplar) (Zhu et al. 2022), and some methods have attempted to learn new classes in a few-shot scenario (Zhang et al. 2021) in which only a small number of new class data is gainable. Attempting to break the one-shot defense assumption, a concurrent work (Zhou and Hua 2024) is proposed to alleviate catastrophic forgetting in a simple continual defense scenario. In this paper, we convert the proposed defense scenario into a few-shot and non-exemplar CL setting, which necessitates the defense mechanism to utilize few-shot defense budget and efficient memory for adaptation.
Threat Model and Defense Principles
@@
The defense system is deployed online and protects the target model against a dynamic environment marked by a plethora of unique attacks. As time progresses, new attacks accompanied by poisoned examples will emerge on the Internet, aiming to exploit vulnerabilities in the target model and achieve breakthroughs. Due to the reality, we propose a practical threat model in which distinct attacks emerge across various stages, requiring the defense system to adapt to evolving threats through the utilization of the defense budget (i.e. adversarial examples and their ground-truths). Typically, a stage can consist of only a few samples from one or multiple attacks, and a new stage begins once the system gathers enough defense budget. The defense budget is crucial as it informs and influences the defender’s response to emerging threats, often discovered during screening or cleanup efforts in the real world, and is provided by web crawlers, security services, and users of the target model. The continual defense scenario is formulated as follows.
@@
Figure 2: Illustration of CAD. Regarding the classes with new attacks as incremental classes, we use continual adaptation to protect the target model while preventing catastrophic forgetting. In the 00-th stage, we reserve the embedding space for future attacks during training of the initial defense model $f_{0}$ to alleviate the overfitting brought by future fine-tuning. Then, a weight estimator model $f_{p}$ is trained for model ensemble. In the $i$-th stage ($i>0$), we freeze the feature extractor $\varphi$ of defense model $f_{i}$ and fine-tune the expanded classifier $g_{c}^{i}$ using few-shot defense budget. To efficiently use memory, we store class-wise prototypes in the prototype cache at each stage and use them in the fine-tuning process. After all, the classifier is refined, and the target model $f_{t}$ and the scalable defense model $f_{i}$ are ensembled by the weight estimator model $f_{p}$ to maintain high performance in classifying clean and adversarial data.
@@
Our scenario is established on the $N$-way $K$-shot classification paradigm. Attacks are launched under the gray-box setting (Taran et al. 2019) in which the attacks know the architecture of the target model ${f}_{t}:\mathcal{X}\to\mathbb{R}^{N}$ which trained on $\mathcal{D}_{\mathrm{train}}$ and evaluated on $\mathcal{D}_{\mathrm{test}}$, but blind to the defense, where $\mathcal{X}$ is the image space. The defender is permitted to visit $\mathcal{D}_{\mathrm{train}}$ and has the full knowledge of the initial attack $A_{0}(\cdot)$ at the 0-th stage. At the $i$-th stage, where $i=1,2,\cdots,T$, new attacks $A_{i}(\cdot)$. are detected and the defender receives a set of budget $\mathcal{A}_{\mathrm{train}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$ containing $N\times K$ samples (i.e., $K$ samples for each of $N$ classes) and uses it to adapt to the new attack. Evaluations are conducted on $\mathcal{D}_{\mathrm{test}}$ and $\{\mathcal{A}_{\mathrm{test}}^{k}\}_{k=0,1,\dots,i}$ at each stage $i$, where $\mathcal{A}_{\mathrm{test}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{test}}\}$.
@@
Continual Adversarial Defense Framework
@@
Continual Adaptation to New Attacks
In response to Principle 1, we use continual adaptation to defend against new attacks while preventing catastrophic forgetting. In the 00-th stage, an initial defense model $f_{0}:\mathcal{X}\to\mathbb{R}^{N}$ which consists of a feature extractor $\varphi:\mathcal{X}\to\mathbb{R}^{d}$ and a classifier $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ is optimized under full supervision using adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}=\{(\mathbf{x}_{\mathrm{adv}}^{0},y)|\mathbf{x}_{\mathrm{adv}}^{0}=A_{0}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$. The defense model is designed to tackle adversarial examples, complementing the target model. In the $i$-th stage, the defense model ${f}_{i}:\mathcal{X}\to\mathbb{R}^{N}$ adapts to the new attack using the defense budget $\mathcal{A}_{\mathrm{train}}^{i}$.
Inspired by continual learning (CL) (Rebuffi et al. 2017), we regard the classes with the new attack as incremental classes and endow the defense model with scalability by classifier expansion. Due to the internal similarity of some adversarial attacks (Wang et al. 2023a), trained with one attack, the defense model can, to some extent, generalize to other attacks, resulting in some unnecessary class expansion. Old attacks occurring in new stages also lead to unnecessary adaptation. Therefore, we conduct model tests and filter out the useless data from the defense budget that can be correctly classified.
Following (Liu et al. 2022), we expand the classifier from $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ to $g_{c}^{1}:\mathbb{R}^{d}\to\mathbb{R}^{N\times{2}}$ in the first stage, and expand the classifier $g_{c}^{i}$ from $\mathbb{R}^{d}\to\mathbb{R}^{N\times i}$ to $\mathbb{R}^{d}\to\mathbb{R}^{N\times{(i+1})}$ in the $i$-th stage. The parameter of the expanded classifier is composed of the parameter of the old classifier and the newly initialized parameter: $W^{i}=[W^{i-1},W^{new}]$, where $W^{i}=[\mathbf{w}^{i}_{1},\cdots,\mathbf{w}^{i}_{N_{i}}]$ is the parameter matrix of $g_{c}^{i}$. Meanwhile, the ground truth of $\mathcal{A}_{\mathrm{train}}^{i}$ is rewritten corresponding to the incremental classes with the new attack  [^2] We omit the networks’ parameter in formulas:

### v5

@@ Abstract
In response to the rapidly evolving nature of adversarial attacks against visual classifiers on a monthly basis, numerous defenses have been proposed to generalize against as many known attacks as possible. However, designing a defense method that generalizes to all types of attacks is not realistic because the environment in which defense systems operate is dynamic and comprises various unique attacks that emerge as time goes on. A well-matched approach to the dynamic environment lies in a defense system that continuously collects adversarial data online to quickly improve itself. Therefore, we put forward a practical defense deployment against a challenging threat model and propose, for the first time, the Continual Adversarial Defense (CAD) framework that adapts to attack sequences under four principles: (1) continual adaptation to new attacks without catastrophic forgetting, (2) few-shot adaptation, (3) memory-efficient adaptation, and (4) high accuracy on both clean and adversarial data. We explore and integrate cutting-edge continual learning, few-shot learning, and ensemble learning techniques to qualify the principles. Extensive experiments validate the effectiveness of our approach against multiple stages of modern adversarial attacks and demonstrate significant improvements over numerous baseline methods. In particular, CAD is capable of quickly adapting with minimal budget and a low cost of defense failure while maintaining good performance against previous attacks. Our research sheds light on a brand-new paradigm for continual defense adaptation against dynamic and evolving attacks.
@@
Figure 1: In a dynamic setting, the target model $f_{t}$ encounters evolving attacks, safeguarded by the defender. At stage $i$, the defense model adapts to new attacks using few-shot defense budget provided by the web crawler, security department, and users of the target model, while efficiently managing memory to retain past attack knowledge.
1 Introduction
@@
Adversarial attack [23] aims to deceive deep neural networks (DNNs) by adding subtle perturbations to input images, seriously jeopardizing the reliability of DNNs, particularly in domains sensitive to security and trust. In order to promote the robustness of DNN, adversarial training [40] and purification techniques [38] are proposed to defend against adversarial attacks by one-shot training in which the model enters a static phase after a single defense training stage [44], but often exhibit reduced classification accuracy in adversarial examples and can sacrifice classification capacity in clean images.
@@
Unlike the one-shot defense assumption, real-world defense systems face dynamic environments with increasingly aggressive and diverse attacks [19]. A well-matched approach to the dynamic environment lies in a constantly evolving defense strategy that continuously adapts to newly emerging adversarial algorithms [3]. However, current model adaptation defenses [2] focus on continuously changing defense strategies when under attack rather than leveraging past defense experiences, leading to suboptimal performance. A concurrent continual defense work [44] treats attacks as a task sequence and tries to train a robust model through lifelong learning. However, poor performance in both adversarial and clean data hinders its deployment in reality and calls into question the reasonableness of the continual defense setting.
@@
To eliminate the insufficiency of the above defense setting, we put forward a practical defense deployment and a challenging threat model. As depicted in Figure 1, the defense system is deployed in the cloud and continuously adapts to new emerging attacks from stage to stage using defense budget (i.e. adversarial examples and their ground-truths) sourced from the web crawler, security department, and users of the target model in real-world scenarios. In reality, a stage can consist of only a few samples from one or multiple attacks, and a new stage begins once the system gathers enough defense budget. When adapted to new emerging attacks, the defense system is permitted to use defense budget of previous stages cached in memory, to maintain the knowledge of previous attacks. The defense system also has access to an initial attack and shares the same training data and architecture as the target model. Meanwhile, attacks are initiated under the gray-box setting, wherein they possess knowledge of the classifier’s architecture and have access to training data, but remain unaware of the defense mechanism implemented.
@@
In this paper, we propose the first Continual Adversarial Defense (CAD) framework, which aims to defend against evolving attacks in a stage-by-stage manner using few-shot defense budget and efficient memory. In the first stage, we use adversarial data from the initial attack to train an initial defense model which is specifically designed to classify adversarial data, complementing the target model. Drawing inspiration from continual learning (CL), we expand the classification layer to adapt to the new attack as incremental classes and then fine-tune this expanded layer using few-shot defense budget. To address the over-fitting issue that arises from the few-shot fine-tuning, we compress the embedding of previous attacks by generating and assigning virtual prototypes to reserve embedding space in the defense model for future attacks. For memory efficiency, we employ prototype augmentation which allows us to maintain the decision boundary of previous stages without the need to store any budget explicitly. After each adaptation, we refine the classifier to avoid unnecessary expansion. Simultaneously, a lightweight model is utilized to ensemble the defense model and the target model by estimating reliable logits for input data, ensuring maximum classification accuracy for both clean and adversarial data.
Extensive experiments on CIFAR-10 and ImageNet-100 show that CAD effectively defends against multi-stage attacks with a few-shot defense budget while maintaining high accuracy on clean data. To the best of our knowledge, this is the first exploration of continual adversarial defense and its real-world deployment.
@@
$\bullet$ To counter evolving adversarial attacks, we put forward a deployment of continual defense strategy with a practical threat model where diverse attacks emerge over stages. The defense system must adapt well to these attacks using a few-shot defense budget and efficient memory, while retaining knowledge of past attacks and ensuring high performance on both clean and adversarial data.
$\bullet$ We propose, for the first time, the Continual Adversarial Defense (CAD) framework that defends against attacks in the dynamic scenario under four practical principles: continual adaptation without catastrophic forgetting, few-shot adaptation, memory-efficient adaptation, and high accuracy in classifying both clean and adversarial data. CAD leverage cutting-edge techniques including continual learning, few-shot learning, non-exemplar class incremental learning, and ensemble learning to satisfy the principles.
@@
2 Related Work
2.1 Adversarial Defense
@@
Researchers have proposed various general defense methods against adversarial attacks [23, 36, 20, 34]. As a standard defense method, adversarial training [41, 40, 8, 35, 17, 5, 11] aims to enhance the robustness of the target model by training it with adversarial examples. Another branch of adversarial defense involves purifying the data stream [26, 10, 38, 24] to remove potential adversarial perturbations or noise that could deceive the model. To counter attacks in dynamic environments, model adaptation methods [2, 12, 37, 18] continuously adjust their parameters, state, or activations when encountering attacks to enhance robustness.
In fact, all of the above methods have shown limitations in robustness, resulting in poor classification performance on both adversarial and clean data. In this paper, we present the first Continual Adversarial Defense (CAD) framework as a solution to effectively counter a wide spectrum of attacks emerging stage by stage, while ensuring robust performance in classifying clean data.
2.2 Continual Learning
@@
Continual learning (CL) [21, 14, 32, 43] aims to learn from a sequence of new classes without forgetting old ones.In recent years, some methods have been aimed at solving the CL problem without relying on preserving data (none-exemplar) [46], and few-shot scenarios [39] in which only a small number of new class data is gainable. Attempting to break the one-shot defense assumption, a concurrent work [44] is proposed to alleviate catastrophic forgetting in a simple continual defense scenario. In this paper, we convert the proposed defense scenario into a few-shot and non-exemplar CL setting, which necessitates the defense mechanism to utilize few-shot defense budget and efficient memory for adaptation.
3 Threat Model and Defense Principles
@@
The defense system is deployed online and shielding the target model from a dynamic environment marked by a plethora of unique attacks. As time progresses, new attacks accompanied by poisoned examples will emerge on the Internet, aiming to exploit vulnerabilities in the target model and achieve breakthroughs. In consideration of this reality, we propose a practical threat model in which distinct attacks emerge over stages, requiring the defense system to adapt to evolving threats through the utilization of the defense budget (i.e. adversarial examples and their ground-truths). Typically, a stage can consist of only a few samples from one or multiple attacks, and a new stage begins once the system gathers enough defense budget. The defense budget is crucial as it informs and influences the defender’s response to emerging threats, often discovered during screening or cleanup efforts in the real world, and is provided by web crawlers, security services, and users of the target model. The continual defense scenario is formulated as follows.
@@
Figure 2: Illustration of CAD. Treating new attack classes as incremental, we use continual adaptation to protect the target model while preventing catastrophic forgetting. In stage 0, we reserve embedding space during initial training to mitigate future overfitting. A weight estimator model $f_{p}$ is trained for model ensemble. In stage $i$ ($i>0$), we freeze the feature extractor $\varphi$ of $f_{i}$, fine-tune the expanded classifier $g_{c}^{i}$ with a few-shot defense budget, and store class-wise prototypes in a cache for efficient memory use. Finally, $f_{t}$ and $f_{i}$ are ensembled via $f_{p}$ for robust classification.
@@
Our scenario is established on the $N$-way $K$-shot classification paradigm. Attacks are launched under the gray-box setting [28] in which the attacks know the architecture of the target model ${f}_{t}:\mathcal{X}\to\mathbb{R}^{N}$ which trained on $\mathcal{D}_{\mathrm{train}}$ and evaluated on $\mathcal{D}_{\mathrm{test}}$, but blind to the defense, where $\mathcal{X}$ is the image space. The defender is permitted to visit $\mathcal{D}_{\mathrm{train}}$ and has the full knowledge of the initial attack $A_{0}(\cdot)$ at stage 0. At stage $i$, where $i=1,2,\cdots,T$, new attacks $A_{i}(\cdot)$ are detected and the defender receives a set of budget $\mathcal{A}_{\mathrm{train}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$ [^1] In the adversarial attack formula $\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x},y,f_{t})$, we omit the target model $f_{t}$ and the ground-truth $y$. containing $N\times K$ samples (i.e., $K$ samples for each of $N$ classes) and uses it to adapt to the new attack, where $y$ is the ground truth of sample $\mathbf{x}$. Evaluations are conducted on $\mathcal{D}_{\mathrm{test}}$ and $\{\mathcal{A}_{\mathrm{test}}^{k}\}_{k=0,1,\dots,i}$ at each stage $i$, where $\mathcal{A}_{\mathrm{test}}^{i}=\{(\mathbf{x}_{\mathrm{adv}}^{i},y)|\mathbf{x}_{\mathrm{adv}}^{i}=A_{i}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{test}}\}$.
@@
4 Continual Adversarial Defense Framework
@@
4.1 Continual Adaptation to New Attacks
In response to Principle 1, we use continual adaptation to defend against new attacks while preventing catastrophic forgetting. At stage 00, an initial defense model $f_{0}:\mathcal{X}\to\mathbb{R}^{N}$ which consists of a feature extractor $\varphi:\mathcal{X}\to\mathbb{R}^{d}$ and a classifier $g_{c}^{0}:\mathbb{R}^{d}\to\mathbb{R}^{N}$ is optimized under full supervision using adversarial dataset $\mathcal{A}_{\mathrm{train}}^{0}=\{(\mathbf{x}_{\mathrm{adv}}^{0},y)|\mathbf{x}_{\mathrm{adv}}^{0}=A_{0}(\mathbf{x}),(\mathbf{x},y)\in\mathcal{D}_{\mathrm{train}}\}$. The defense model is designed to tackle adversarial examples, complementing the target model. At stage $i$, the defense model ${f}_{i}:\mathcal{X}\to\mathbb{R}^{N}$ adapts to new attacks using the defense budget $\mathcal{A}_{\mathrm{train}}^{i}$ while mitigating catastrophic forgetting, as outlined in Principle 1, with cached prototypes $P_{e}$ extracted from past attacks.
Inspired by continual learning (CL) [25], we regard the classes with the new attack as incremental classes and endow the defense model with scalability by classifier expansion. Following [21], in stage $i$ $(i>0)$, we expand the classifier $g_{c}^{i-1}:\mathbb{R}^{d}\to\mathbb{R}^{N\times i}$ to $g_{c}^{i}:\mathbb{R}^{d}\to\mathbb{R}^{N\times{(i+1})}$. The parameters of the expanded classifier consist of those from the old classifier and newly initialized ones: $W^{i}=[W^{i-1},W^{new}]$, where $W^{i}=[\mathbf{w}^{i}_{1},\cdots,\mathbf{w}^{i}_{N_{i}}]$ is the parameter matrix of $g_{c}^{i}$. Meanwhile, the ground truth of $\mathbf{x}_{\mathrm{adv}}^{i}\in\mathcal{A}_{\mathrm{train}}^{i}$ is rewritten corresponding to the incremental classes with the new attack  [^2] We omit the networks’ parameter in formulas.: